MiraAngel:

Given 2x^3-15x^2+27x-10=0, where x=1/2, use synthetic division to show that x is a solution of the third-degree polynomial equation and use the result to factor the polynomial completely. List all real solutions.

MiraAngel:

There are no diagrams to upload for this problem. I got a messages on here to make sure I upload any. Just wanted to let anyone who comes this post that there's no diagram to go with it

MiraAngel:

@Isolation

Isolation:

I'm confused... it gives you the x-value.... but what's you to use synthetic division? to divide what? or does it mean to divide all that with (x-1/2)

Isolation:

If that's the case... I'd prefer to use long division... lol

Isolation:

But you still can use long division... but 1/2 as that little number thing what'd you get if you do it

jhonyy9:

so given x = 1/2 or 2x = 1 from what result that you can divide this equation by (x - 1/2) or (2x - 1)

surjithayer:

do you know the process of synthetic division? write coefficients of the given eq. and then start. 1/2|2 -15 27 -10 write 2,then multiply by 1/2,you get 1 ,write and add you get -14 | 1 -7 10 again multiply by 1/2,you get -7,write and add ,you get 20 , _________________ again multiply by 1/2,you get 10 ,again multiply by 1/2 ,you get 10, 2 -14 20 | 0 add ,remainder =0 now you get aquadratic eq. can you write it?

surjithayer:

i have written twice agai multiply by 1/2,you get 10 by mistake

surjithayer:

it is only once.

surjithayer:

can you make factors of this quadratic?

MiraAngel:

O_O that's a lot to take in lol